OpenStudy (anonymous):

need help please 5 3/4 - 2 1/2

OpenStudy (anonymous):

\[5\frac{ 3 }{ 4 } - 2\frac{ 1 }{ 2 }\]

OpenStudy (anonymous):

is this correct...

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok then the convert the mixed numbers into improper fractions like so: 23/4 - 5/2 take the LCM and then solve the fractions

OpenStudy (anonymous):

sorry its 5 3/4 - 2 1/2

OpenStudy (anonymous):

\[\frac{ 53 }{ 4 } - \frac{ 21 }{ 2 }\]

OpenStudy (anonymous):

is that correct????

OpenStudy (anonymous):

how did you get that??

OpenStudy (anonymous):

no im asking u the question...what is the question... write it in the equation form...

OpenStudy (anonymous):

\[5\frac{ 3 }{ 4} - 2 \frac{ 1 }{ 5 } \]

OpenStudy (anonymous):

to the lowest term

OpenStudy (paki):

23/4-11/5... is it right one @jovi808

OpenStudy (anonymous):

what do i do nexts?

OpenStudy (paki):

ok... take the LCM.... then 23/4-11/5 115-110/20.... lcm is 20 here ok... 5/20 1/4

OpenStudy (paki):

@jovi808 got it...

OpenStudy (anonymous):

where did the 110 come from??

OpenStudy (paki):

oh sorry... its 44.. :-( 115-44/20= 71/20.... this is rite now.... actually i am solving 2-3 questions at the same time here, so, that's why error... :_)

OpenStudy (anonymous):

got it thanks

OpenStudy (paki):

pleasure...

OpenStudy (anonymous):

\[1\frac{ 5 }{ 12 } + 2\frac{ 7 }{ 9 }\]

OpenStudy (paki):

26/12+25/9 take lcm... which is 36 here 15+28/36 43/36 1.96

OpenStudy (paki):

@jovi808

OpenStudy (anonymous):

so how did u get 26/12 +25/9??

OpenStudy (paki):

again the same problem.... sorry... its 17/12+25/9 :-(

OpenStudy (anonymous):

got that

OpenStudy (anonymous):

what do i do next?

OpenStudy (paki):

take the lcm... and solve

OpenStudy (paki):

51+100/36 = 151/36 = 4.19

OpenStudy (anonymous):

how did you get that??

OpenStudy (paki):

what

OpenStudy (anonymous):

how did u get the 51 n 100??

OpenStudy (paki):

36/12 * 17 = 51

OpenStudy (paki):

36/9 * 25 = 100

OpenStudy (paki):

@jovi808 got it now :-)

OpenStudy (anonymous):

yes

OpenStudy (paki):

thanks :-)